Ask
your own question, for FREE!
Algebra
10 Online
You toss 3 coins. Find P(at least 1 tail is tossed).
Still Need Help?
Join the QuestionCove community and study together with friends!
75%
1 tail is 50%
So 75% ?
yea
for three
Still Need Help?
Join the QuestionCove community and study together with friends!
for three tosses it is 75%
Thank you sorry im just so tierd its hard to think.. ;(
np yw
87%??
ok cool
Still Need Help?
Join the QuestionCove community and study together with friends!
P(at least 1 tail is tossed)= 1 - p(no tail at all) p(no tail )means all three heads p(all three heads ) =1/8 hence P(at least 1 tail is tossed)= 1- 1/8 =7/8
It was 7/8 ! Lol
:/
I didn't get it matrick
-ed
Still Need Help?
Join the QuestionCove community and study together with friends!
when you toss a coin it is either heads or tails... getting a tails is 1/2 which is 50% right??? so half of 150% ( three times) is what?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal